OpenStudy (anonymous):

Luis is in charge of making props for a school play. He needs to make a large circular wooden clock that measures about 6 feet in circumference. Which equation can he use to find r, the radius of the clock? A. r = 6/pi B. r= 12/pi C. r=6/2pi D. r= 12/2pi

OpenStudy (anonymous):

What is the equation for the circumference of a circle?

Parth (parthkohli):

2 * pi * radius

OpenStudy (anonymous):

C= pi times diameter

OpenStudy (anonymous):

But what does diameter equal in terms of radius?

Parth (parthkohli):

@Romero radius = diameter/2

OpenStudy (anonymous):

half of it.

OpenStudy (anonymous):

the radius is half the diameter

OpenStudy (anonymous):

the formula is, 2pi(r) = C

OpenStudy (anonymous):

would the answer be C?

OpenStudy (anonymous):

radius = diameter/2 or 2* radius= diameter so lets replace this with the equation of the circumference C= pi *2* radius And we know the circumference is 6 so lets replace that 6= pi *2* radius

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

If you rearrange to get r isolated then yes the answer will be c

OpenStudy (anonymous):

@Hero I got this hero :P

hero (hero):

1. Wriite out the formula for circumference: \[C = 2 \pi r\] 2. Input the given value for the circumference, 6 into the formula: \[6 = 2 \pi r\] 3. Divide both sides by 2 pi: \[\frac{6}{2\pi} = r\] 4. Compute r:

hero (hero):

@Romero , if you really had it, then I wouldn't have to type anything.

OpenStudy (anonymous):

, THANKS.

OpenStudy (anonymous):

@Hero I dud but you interrupted with the answer. You know that's not allowed right?

OpenStudy (anonymous):

So, final answer is C,right?

hero (hero):

I provided steps and explained every step. I didn't do anything wrong.

OpenStudy (anonymous):

its C 6/2pi

Parth (parthkohli):

Oops, I deleted my answer by mistake :(

OpenStudy (anonymous):

ok, thanks, everyone.

OpenStudy (anonymous):

@Hero No you can't interrupt someone who is already explaining it. It just confuses the asker.

Parth (parthkohli):

@Romero agree

hero (hero):

Romero, I interrupted because your explanations did not seem clear.

hero (hero):

@happyhorse can probably back me up on that.

OpenStudy (anonymous):

Yes, hero.

OpenStudy (anonymous):

@Hero It wasn't clear because you interrupted

OpenStudy (anonymous):

He didnt confuse me at all

Parth (parthkohli):

We can calculate by tranposing. Here it is :